OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. 1. 1 + sec^2x sin^2x = sec^2x 2. (sinx / 1 - cosx) + (sinx / 1 - cosx) = 2 csc x 3. - tan^2x + sec^2x = 1 It would be great to be shown step by step how to solve these problems. Thanks in advance!

OpenStudy (anonymous):

For number one, I believe the answer is: 1 + sec^2(x) sin^2(x) = sec^2(x) sec(x) = 1/cos(x) sec(x)sin(x) tan(x) 1 + tan^2(x) sec^2(x), which equals the right side. Please correct me if I'm wrong!!

OpenStudy (ipwnbunnies):

Well, almost correct. sec^2(x)*sin^2(x) = tan^2(x) But you had the correct process.

OpenStudy (anonymous):

Ok, thanks! And for the second question I got: (sin x) / (1 - cos x) + (sin x) / (1 + cos x) = 2 csc x (1 - cos x)(1 + cos x) (each side is then multiplied with this) (sin x)(1+cos x) + (sin x)(1 - cos x) = 2 csc x (1 - cos^2 x) 1 - cos^2 x = sin^2 x (sin x)(1 + cos x) + (sin x)(1 - cos x) = 2 csc x (sin^2 x) sin x + sin x cos x + sin x - sin x cos x = 2 sin x 2 sin x I think I did this right, because 1 / sinx equals cscx, correct?

OpenStudy (anonymous):

You are very correct for doing that..

OpenStudy (anonymous):

Yep, Good..

OpenStudy (anonymous):

Wait, how you got 2isn(x)

OpenStudy (anonymous):

*2sin(x)

OpenStudy (anonymous):

Hm, Idk I just solved..lol is it wrong?

OpenStudy (anonymous):

Oh ok, thanks

OpenStudy (anonymous):

Sorry got you.. :)

OpenStudy (anonymous):

So I'm right? :P

OpenStudy (anonymous):

\[\frac{2\sin(x)}{\sin^2(x)} \implies \frac{2}{\sin(x)} \implies 2 cosec(x)\]

OpenStudy (anonymous):

I see, thx

OpenStudy (anonymous):

I think third one you can do it, right??

OpenStudy (anonymous):

I can, I haven't done it yet though. I'll have you check my answer when I'm done, ok? :)

OpenStudy (anonymous):

For third one, you have to do anything??

OpenStudy (anonymous):

In first equation, you used: \(1 + tan^2(x) = sec^2(x)\) Just rearrange it by subtracting \(tan^2(x(\) from both the sides, you will get your desired thing..

OpenStudy (anonymous):

So it's just 1?

OpenStudy (anonymous):

No, its 1 = sec^2(x) - tan^2(x)

OpenStudy (anonymous):

Oops, that's what I meant..forgot to subtract tan^2x from sec^2x lol

OpenStudy (anonymous):

And that's the answer, right? :) Thanks for your help!

OpenStudy (anonymous):

yes.. :)